OpenStudy (anonymous):

Determine the number that will complete the square to solve the equation after the constant has been written on the right-hand side.

OpenStudy (anonymous):

\[x^{2} -14x-15=0\]

OpenStudy (anonymous):

Remember that if you had the following square: \[ (a+b)^2 \]And expand it you get: \[ (a+b)^2 = a^2 + 2ab + b^2 \]So to complete the square, you need to identify \(a\) and \(b\). Then you need to pull out \(b\).

OpenStudy (anonymous):

We start by identifying \(a=x\).

OpenStudy (anonymous):

This means: \(2xb = -14x\). So you solve for \(b\). Can you do it? @cralls18

OpenStudy (anonymous):

Basically \[ (a+b)^2 = a^2 + 2ab + b^2 \\ \]With \[ x^2 - 14x-15 \]We have \(a = x\). This means that: \[ 2ab = -14x \implies 2xb = -14x \]

OpenStudy (anonymous):

@cralls18 Can you solve for \(b\) here?

OpenStudy (anonymous):

do u subtract 2x? from both sides.

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

so -16x?

OpenStudy (anonymous):

Whoops... I misread. You don't subtract, you divide both sides by \(2x\).

OpenStudy (anonymous):

-7?

OpenStudy (anonymous):

Yes. So our square is \[ (x+(-7))^2 \]Can you expand it using FOIL?

OpenStudy (anonymous):

ok.. x(sqrd)-14x+49

OpenStudy (anonymous):

Yes. \[ (x-7)^2 = s^2-14x+49 \]But the problem is that this isn't equal to: \[ x^2 -14x -15 \]

OpenStudy (anonymous):

not \(s\) meand \(x\) up there.

OpenStudy (anonymous):

A. -4 B. 10 C. 49 D. None of these. E. 4 those are my answers.....

OpenStudy (anonymous):

We want to find: \[ (x-7)^2 +c = x^2 -14x-15 \]Expanding this gets: \[ x^2 - 14x + 49 + c = x^2 -14x-15 \]Can you solve for \(c\)?

OpenStudy (anonymous):

Actually, maybe we've done enough.

OpenStudy (anonymous):

Just take the coefficient of x (-14) divide it by 2, and then square that number... (-14/2)^2

OpenStudy (anonymous):

It looks like they're asking for 49.

OpenStudy (anonymous):

ok. thx